OpenStudy (lilai3):

please help for a riddle! A store is offering a $30 mail-in rebate on all color printers. Luisana is looking at different color printers that range in price from $175 to $260. How much can she expect to spend after the mail-in rebate? please explain. thank you.

OpenStudy (alyssajobug):

175-260, she has to mail in the rebate with proof of purchase before she can get reenbursed

OpenStudy (lilai3):

so... 175-30 or 260-30, then?

OpenStudy (mathstudent55):

correct $175 - $30 to $260 - $30

OpenStudy (lilai3):

that is the inequality?

OpenStudy (mathstudent55):

If she spends $175, then after the $30 rebate she will have spent $145. If she spends $260, then after the $30 rebate she will have spent $230. That means she can spend between $175 and $230, inclusive. \( 145 \le x \le 230\)

OpenStudy (lilai3):

why the <= sign though?

OpenStudy (mathstudent55):

You need <= because the 175 and the 230 values are included.

OpenStudy (lilai3):

oh i get it. thank you very much @mathstudent55

OpenStudy (mathstudent55):

wlcm

OpenStudy (lilai3):

ahhh have uno de problemo. please help/ The recommended air pressure for the tires of a mountain bike is at least 35 pounds per square inch (psi), but no more than 80 pounds per square inch. If a bike's tires have 24 pounds per square inch, what increase in air pressure is needed so the tires are in the recommended range?

OpenStudy (lilai3):

un problema***

OpenStudy (lilai3):

anyway, it's alright. you already left. lool here's your riddle: What has ten letters and starts with gas?

OpenStudy (lilai3):

solve it, and then when you give up... scroll d o w n . . . . .

OpenStudy (lilai3):

.

OpenStudy (lilai3):

.

OpenStudy (lilai3):

.

OpenStudy (lilai3):

.

OpenStudy (lilai3):

.

OpenStudy (lilai3):

.

OpenStudy (lilai3):

.

OpenStudy (lilai3):

.

OpenStudy (lilai3):

.

OpenStudy (lilai3):

an automobile! ;)

OpenStudy (mathstudent55):

The tire has 24 psi. It needs at least 35 psi to have the proper pressure. 35 psi - 24 psi = 11 psi. If you increase the pressure by at least 11 psi, the pressure will be already 35 psi or higher. Now you need to figure out what is the most you can increase the pressure by. Since the highest pressure allowed is 80 psi, then do 80 psi - 24 psi = 56 psi. The most you can increase the pressure and still be ok is 56 psi. The final inequality is (using p for pressure): \( 11 ~psi \le p \le 56 ~psi\)

OpenStudy (lilai3):

oh i see. thank you so much for your help again!

OpenStudy (mathstudent55):

You're welcome.